Solve: 12(3z + c) - 2(2z + 4c)

Solve: 3x-4=2x + 8

Answer by user_dude2008(1862)

12(3z + c) - 2(2z + 4c) = 36z+ 12c - 4z -8c
........................= 32z + 4c

Ans: 12(3z + c) - 2(2z + 4c) = 32z + 4c





3x-4=2x + 8

3x-2x = 8 + 4

x = 12

Ans: x = 12